OpenStudy (theopenstudyowl):

If 1+2 = 13, 2+5 = 37, 21 - 41 = -2020, 76 - 40 = 3636, 10 + 1 = 911, what does 2 + 6 = ?

OpenStudy (afrodiddle):

42 Google says that is the answer to everything. Is this question fore real, because I am completely lost.

OpenStudy (anonymous):

48

OpenStudy (theopenstudyowl):

haha, 2+6 = 48; the difference between the two number is the first number, and then the sum or difference is written next..... the difference between 2 and 6 is 4, and the sum is 8! very good @lokeshh

OpenStudy (afrodiddle):

~lost.

Nnesha (nnesha):

-.-

OpenStudy (theopenstudyowl):

does it make sense?

OpenStudy (afrodiddle):

Not at all :S

OpenStudy (theopenstudyowl):

ok what is 2+3 = 15

Nnesha (nnesha):

wrong

Nnesha (nnesha):

2-3 = -1 :o

OpenStudy (afrodiddle):

-15

Nnesha (nnesha):

^

OpenStudy (theopenstudyowl):

the difference is 1, and the sum of 2 and 3 is 5

OpenStudy (theopenstudyowl):

@Nnesha , I know its not right, but for the sake of the logic test, that is our function/way of thinking if u will..

Nnesha (nnesha):

2+1 = 3 2-1 = 1 2+1 =13

Nnesha (nnesha):

LOL

OpenStudy (theopenstudyowl):

lol

OpenStudy (afrodiddle):

Are you preparing for an IQ test or something? o.o

OpenStudy (theopenstudyowl):

haha no, but I enjoy questions like this!

Nnesha (nnesha):

wrong :p 2+5 = -37 5+2 = 37 o^_^o

OpenStudy (afrodiddle):

Find an IQ test online :P They will give you a butt load of logic questions.

OpenStudy (theopenstudyowl):

lol

OpenStudy (theopenstudyowl):

yah! for sure, gotta get back to work though lol

OpenStudy (theopenstudyowl):

Make an account and join Openstudy! It's free!!!!!

Nnesha (nnesha):

me ?

OpenStudy (theopenstudyowl):

no @Nnesha there was a person who was looking at my question from outside the website.....